# What Determines the Maximum Mass of AGN-assisted Black Hole Mergers?

### v1

@@ Abstract
The origin of merging binary black holes detected through gravitational waves remains a fundamental question in astrophysics. While stellar evolution imposes an upper mass limit of $\sim 50$Mdirect-$\odot$ for black holes, some observed mergers—most notably GW190521—involve significantly more massive components, suggesting alternative formation channels. Here we investigate how heavy black holes merging within Active Galactic Nucleus (AGN) disks can become. Using a comprehensive semi-analytic model incorporating 27 binary and environmental parameters, we explore the role of AGN disk conditions in shaping the upper end of the black hole mass spectrum. We find that AGN disk lifetime is the dominant factor, with high-mass mergers ($\gtrsim 200$Mdirect-$\odot$) only possible if disks persist for $\gtrsim 40$ Myr. The joint electromagnetic observation of an AGN-assisted merger could therefore lead to a direct measurement of the age of an AGN disk.
@@ 1 Introduction
Binary black hole mergers are now regularly being observed through gravitational wave emission. The LIGO (Aasi et al., 2015), Virgo (Acernese et al., 2015) and KAGRA (Akutsu et al., 2019) collaborations have reported around 90 previous discoveries, while their 4th observing run, O4, is currently ongoing.
@@
The high end of the observed black hole mass distribution is a particularly promising part of the parameter space. Generally, heavier black holes are expected to be rarer, making processes that preferentially select heavier black holes for mergers stand out. Even more importantly, black hole formation from stellar evolution is expected to reach only up to a mass of around 505050\,50M (Fryer et al., 2001). Massive stars above a critical mass are expected to experience pair instability in their center, resulting in their early explosion before they could form a sufficiently massive core. These stars are expected to leave no compact remnant behind, effectively putting an upper limit on how massive black holes can be formed from stars.
Black holes have nonetheless been observed with over 505050\,50M mass. In particular, merger GW190521 was recorded to have been a binary of two black holes with masses estimated to be $95.3^{+28.7}_{-18.9}$ and $69.2^{+17.0}_{-10.6}$ M (Abbott et al., 2020, 2021). Such high masses are difficult to explain with stellar evolution (although see Belczynski 2020). In addition, GW190521 featured other outlying properties, including its apparent high precessing spin, and possibly highly eccentric orbit prior to merger (Gayathri et al., 2022; Romero-Shaw et al., 2020; Gamba et al., 2023), pointing towards a dynamical or AGN origin (Samsing et al., 2022).
A promising mechanism to explain the creation of black hole masses beyond stellar evolution limits is the consecutive merger of black holes, or so-called hierarchical mergers (Gerosa & Berti, 2017; Fishbach et al., 2017; Yang et al., 2019; Tagawa et al., 2020b, 2021; Mahapatra et al., 2021, 2024, 2025). The remnant of the merger of two black holes can naturally exceed the maximum mass allowed by stellar evolution, therefore, if such remnants can participate in further mergers, possibly much higher black hole masses can be reached.
@@
In this paper we explore this connection between maximum mass and environment. The goal of this investigation is to understand how the observation of the heaviest black holes can be converted to understanding of the environment of the black holes, or astrophysical processes more generally.
We will in particular focus on black hole mergers in AGN disks. AGNs are promising black hole assembly lines that are expected to promote hierarchical mergers (Yang et al., 2019), while also may grow black hole masses through accretion (Yang et al., 2020).
@@
The paper is organized as follows. Section 2 presents our AGN model and binary gas capture, evolution and merger simulation framework. We present our results in Section 3. We discuss some of the limitations of our model and future work in 4, and conclude in Section 5.
@@ 2 Method
Our AGN model largely builds on the framework established by Tagawa et al. (2020b). Our main modification compared to this model aims to enable faster simulations to be able to cover a larger part of the parameter space. Instead of employing N-body simulations, we adopt a statistical approach to efficiently calculate background properties, significantly reducing computational costs. Especially, this approach enables us to simulate systems with high mass SMBHs and large populations of stellar-mass BHs. Comparable studies using statistical methods have been conducted, focusing on dynamical interactions (Gondán et al., 2018) and gas interactions (McKernan et al., 2012; Vaccaro et al., 2024; Gayathri et al., 2021; Yang et al., 2019) independently. Additionally, Fokker-Planck numerical methods have been used to investigate extreme mass ratio inspirals near SMBHs (Pan & Yang, 2021). Many studies also explore spin evolution due to these mechanisms (Gondán et al., 2018; Vaccaro et al., 2024; Tagawa et al., 2020a; Gayathri et al., 2021). In the present paper, we consider both the gas hardening process and star interactions, while spin evolution will be explored in a future paper.
The BH and stellar populations are described in Section 2.1, which provides their spatial and mass distribution. The Keplerian speed is slightly modified by the presence of the spatially extended stellar populations (Eq. 6).
The properties of the AGN disk are calculated using the model described in Section 2.2. The inner ($r_{\mathrm{disk,in}}$) and outer ($r_{\mathrm{disk,out}}$) boundaries of the AGN disk are treated as free parameters to estimate their effect, though it could be set by the innermost stable circular orbit (ISCO) of SMBH. Details are discussed in section 2.2 and 2.4.3. The simulation begins by determining the accretion rate at the outer boundary of the AGN disk, $\dot{M}_{\mathrm{out}}$, and uses this to compute the disk’s properties from the outer radius inward. It is assumed that stars and BHs are formed in the outer region of the AGN disk, which helps to stabilize the disk. Star formation ceases when the accretion heat at the outer boundary is sufficient to prevent gravitational collapse. This typically occurs in systems with either a low SMBH mass or a low accretion rate at the outer boundary. Given the non-zero star formation rate, disk properties are calculated on a denser radial grid, as suggested in Section 2.2, and then translated to the standard values at grid boundaries and center of $N_{\mathrm{cell}}=120$.
[…]
With the background fully established, we begin by considering single or binary BHs, characterized by BH mass $M_{\mathrm{BH}}$, radial position $r$ from the SMBH, velocity components $v_{x,y,z}$ relative to the disk velocity, and, for binaries, the mass ratio $q$ and separation $s$. We examine the evolution of these quantities as the BHs interact gravitationally (Section 2.3), with the surrounding gas (Section 2.4) and with the surrounding stellar objects as an evolving background (Section 2.5).
Figure 2 illustrates all interaction types considered in the program. In the gravitational interaction, we discuss binary hardening due to gravitational wave emission (Section 2.3.1) and, eventually, the merger into a new single BH (Section 2.3.2). The interaction between the BH and the surrounding gas affects mass, position, and velocity through gas accretion (Section 2.4.1), gas dynamical friction (Section 2.4.2), and Type I/II migration (Section 2.4.3). Gas dynamical friction may also assist in binary formation (Section 2.5.4) and binary hardening process (Section 2.4.4). The interaction between the BH or BH binary and the evolving background includes both weak encounters (weak scattering, Section 2.5.1) and strong encounters, such as binary-single interactions (Section 2.5.2) and binary formation through three-body interactions (Section 2.5.3). These effects renew the BH sample’s properties iteratively with time step given in Section 2.7 until the AGN disk dissipates.
We record all merger events that occur during the simulation, aiming for at least 10,000 mergers. However, since the merger rate varies significantly, strictly requiring 10,000 mergers may lead to excessive computational time. To balance efficiency and accuracy, we maintain a minimum of 10 AGNs if the merger rate is low, where the initial number of BHs $N_{\mathrm{BH},\mathrm{ini}}$ is set by default according to Eq. 9. In the individual sample simulation, the radial distance $r$, Keplerian speed $v_{\mathrm{kep}}$ and the disk properties (e.g. gas density $\rho$) are represented by the cell center values to reduce computational cost.
Our program is written in C++ to optimize computational performance. The AGN disk module takes approximately 7 seconds per AGN and shows minimal variation across parameter choices, as the runtime is primarily dominated by data input/output operations. The statistical background calculation for the fiducial models (both Gaussian and isotropic) requires roughly 600 seconds (CPU time) and scales approximately linearly with the AGN lifetime and the number of time intervals specified in Eq. 113. This computational cost also decreases with increasing SMBH mass.
@@
AGN disk $r_{\mathrm{disk},\mathrm{in}}$ Inner bound of AGN disk $10^{-4}$ pc, following Eq. 29 $0.5\times 10^{-4}$ $1.5\times 10^{-4}$ pc
AGN disk $r_{\mathrm{disk},\mathrm{out}}$ Outer bound of AGN disk $5$ pc, following Eq. 30 5 - 10 pc
AGN disk $\dot{M}_{\mathrm{out}}$ Accretion rate at the outer bound of AGN disk $0.1\dot{M}_{\mathrm{Edd}}$ $0.025\dot{M}_{\mathrm{Edd}}$ $0.20\dot{M}_{\mathrm{Edd}}$
@@
Mass distribution $\beta_{\mathrm{IMF}}$ IMF index of pre-existing BH 2.35 1.7 - 2.35
@@
Radial distribution $N_{\mathrm{\mathrm{BH},\mathrm{ini}}}$ Initial BH number in the AGN (within $r_{BH,out}$) 20000, following Eq. 9 20000 - 100000
Radial distribution $r_{\mathrm{BH},\mathrm{out}}$ Maximum BH radial distance to SMBH of pre-existing BH following Eq. 3 2.5 - 3.5 pc
@@
Binary $f_{\rm{pre}}$ Pre-existing binary BH fraction 0.15 0 - 0.2
[…]

### v2

@@ Abstract
[…]
@@ 1 Introduction
Binary black hole mergers are now regularly being observed through gravitational wave emission. The LIGO (Aasi et al., 2015), Virgo (Acernese et al., 2015) and KAGRA (Akutsu et al., 2019) collaborations have reported around 90 previous discoveries, while their fourth observing run, O4, is currently ongoing.
@@
The high end of the observed black hole mass distribution is a particularly promising part of the parameter space. Generally, heavier black holes are expected to be rarer, making processes that preferentially select heavier black holes for mergers stand out. Even more importantly, black hole formation from stellar evolution is expected to reach only up to a mass of around 5050\,M (Fryer et al., 2001). Massive stars above a critical mass are expected to experience pair instability in their center, resulting in their early explosion before they could form a sufficiently massive core. These stars are expected to leave no compact remnant behind, effectively putting an upper limit on how massive black holes can be formed from stars.
Black holes have nonetheless been observed with over 5050\,M mass. In particular, merger GW190521 was recorded to have been a binary of two black holes with masses estimated to be $95.3^{+28.7}_{-18.9}$M and $69.2^{+17.0}_{-10.6}$ M (Abbott et al., 2020, 2021). Such high masses are difficult to explain with stellar evolution (although see Belczynski 2020). In addition, GW190521 featured other outlying properties, including its apparent high precessing spin, and possibly highly eccentric orbit prior to merger (Gayathri et al., 2022; Romero-Shaw et al., 2020; Gamba et al., 2023), pointing toward a dynamical or AGN origin (Samsing et al., 2022).
A promising mechanism to explain the creation of black hole masses beyond stellar evolution limits is the consecutive merger of black holes, or so-called hierarchical mergers (Gerosa & Berti, 2017; Fishbach et al., 2017; Yang et al., 2019; Tagawa et al., 2020b, 2021; Mahapatra et al., 2021, 2024, 2025). The remnant of the merger of two black holes can naturally exceed the maximum mass allowed by stellar evolution; therefore, if such remnants can participate in further mergers, possibly much higher black hole masses can be reached.
@@
In this paper we explore this connection between maximum mass and environment. The goal of this investigation is to understand how the observation of the heaviest black holes can be converted to understanding of the environment of the black holes, or astrophysical processes more generally. While AGN disks are the primary focus of this work, it is important to note that other environments may also facilitate the formation of intermediate-mass black holes (IMBHs). For example, globular clusters and nuclear star clusters (Chattopadhyay et al., 2023; Barber & Antonini, 2025; Purohit et al., 2024; Atallah et al., 2023; Askar et al., 2017; Gieles et al., 2025), and possibly gas-rich dense stellar systems can support repeated mergers or gas accretion.
We focus in particular on black hole mergers in AGN disks. AGNs are promising black hole assembly lines that are expected to promote hierarchical mergers (Yang et al., 2019), and also may grow black hole masses through accretion (Yang et al., 2020).
@@
The paper is organized as follows. Sec. 2 presents our AGN model and binary gas capture, evolution and merger simulation framework with additional technical details provided in the Appendices A-G . We present our results in Sec. 3. We discuss some of the limitations of our model and future work in 4, and conclude in Sec. 5.
@@ 2 Method
AGN disk $r_{\mathrm{disk},\mathrm{in}}$ inner bound of AGN disk $10^{-4}$ pc, following Eq. B14 $0.5\times 10^{-4}$-$1.5\times 10^{-4}$ pc
AGN disk $r_{\mathrm{disk},\mathrm{out}}$ outer bound of AGN disk 55 pc, following Eq. B15 5-10 pc
AGN disk $\dot{M}_{\mathrm{out}}$ accretion rate at the outer bound of AGN disk $1\dot{M}_{\mathrm{Edd}} 0.0$$025\dot{M}_{\mathrm{Edd}}-$$0.20\dot{M}_{\mathrm{Edd}}$
@@
Mass distribution $\beta_{\mathrm{IMF}}$ IMF index of preexisting BH 2.35 1.7-2.35
@@
Radial distribution $N_{\mathrm{\mathrm{BH},\mathrm{ini}}}$ initial BH number in the AGN (within $r_{BH,out}$) 20000, following Eq. A9 20000-100000
Radial distribution $r_{\mathrm{BH},\mathrm{out}}$ maximum BH radial distance to SMBH of preexisting BH following Eq. A3 2.5-3.5 pc
@@
Binary $f_{\rm{pre}}$ preexisting binary BH fraction 0.15 0-0.2
Binary $R_{\max}$ maximum separation of preexisting binary BH $10^{5}R_{\odot}$ $10^{4}\;-\;10^{5}\,R_{\odot}$